# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ 2.2 SED Fitting
We determine the spectral energy distribution (SED) for each CECILIA target using the archival photometry described above. Note that the Palomar/WIRC photometry in is generally of worse quality than other bands, but in the case of non-detections in the IRAC channels, it is required to constrain the SED redward of the Balmer break. As a result, we only include the photometric point in the fit if the galaxy is undetected in IRAC 1, 2, and 3.
@@
We use BAGPIPES (A. C. Carnall et al., 2018) for SED fitting, using similar methods to those described by N. A. Korhonen Cuestas et al. (2025). We employ the Bayesian nested sampler Nautilus (J. U. Lange, 2023) and the BPASS v.2.2.1 (E. R. Stanway & J. J. Eldridge, 2018) stellar grids assuming a P. Kroupa (2001)-like IMF with upper mass slope of 2.35 and a maximum stellar mass of 100 . Cloudy (G. J. Ferland et al., 2017) grids are used to model the nebular continuum. We do not include nebular emission lines in the SED model, as the photometry has been corrected for the contributions of emission lines using the measured JWST/NIRSpec equivalent widths as well as previously-measured emission line fluxes from MOSFIRE where available. We adopt uniform priors on the dust extinction, , logarithm of the ionization parameter, stellar metallicity, and logarithm of the stellar mass, allowing each parameter to vary between , , , and respectively. Gas phase metallicity is fixed to the stellar metallicity.111Note that the SED fits are insensitive to gas-phase metallicity because the emission lines are subtracted from the photometry and are not fit with the BAGPIPES modeling. Dust extinction is modeled assuming the K. D. Gordon et al. (2003) SMC dust reddening curve.
@@
The SFH is fit using a non-parametric, continuity SFH, as described in J. Leja et al. (2019). We define time bins 0, 10, 100, 250, 500, 1000, 1500, 2000, 2500, 3000, and 3500 Myr before the time of observation and fit for the change in SFR between each bin. In practice, the last time bin is sometimes excluded by the redshift of the galaxy. The prior for changes in SFR between adjacent bins is set to the BAGPIPES default for a continuity SFH; i.e., a Student’s t distribution centered at 0 with a scale and 2 degrees of freedom. Such a prior is weighted against dramatic changes in SFR between time bins, generally producing smoother SFHs. Notably, the nominal SFR reported by BAGPIPES is measured on 100 Myr timescales by default, which produces an apparent maximum specific star formation rate (sSFR) of and artificial linear features to the distribution of objects in the SFR-M∗ plane. However, the underlying SFH used for the SED fitting and total mass estimation includes the full set of time bins described above. The resulting distribution of SED-fit SFRs and stellar masses for the CECILIA sample and their parent population are discussed further in Sec. 2.8.
@@ 2.3 Keck/MOSFIRE Spectroscopy
The Keck/MOSFIRE observations of KBSS galaxies have been described extensively in past work (C. C. Steidel et al., 2014; A. L. Strom et al., 2017). In brief, KBSS used MOSFIRE’s slit configurable slit unit (CSU) to observe galaxies at a time in a single atmospheric window; individual galaxies have observations spanning the , , and bands, which collectively cover Å for our highest-priority targets. Individual mask configurations were designed with -wide slits, resulting in a spectral resolution of . Observations were acquired using a two-position dither sequence along the slit, using the standard dither spacing of . Individual integration times were 180 s in band and 120 s in the and bands; total integration times range from hr/band, with a typical total integration time of hr/band. In general, once a spectroscopic redshift was determined, galaxies continued to be observed until the strongest emission lines (including [O II], H, [O III], H, and [N II]) were detected at , so some of the fainter galaxies in the sample have substantially longer exposure times.
@@
1D spectra are extracted using MOSPEC,333https://github.com/allisonstrom/mospec an interactive IDL-based analysis tool developed for faint emission line spectroscopy with MOSFIRE. We adopt a Gaussian weight profile and optimally extract the spectra using the approach from K. Horne (1986). To account for slit losses, which can differ from band to band depending on observing conditions, we use the Markov Chain Monte Carlo (MCMC) method from A. L. Strom et al. (2017). First, we compare observations of stars that are observed on each MOSFIRE mask with their NIR magnitudes. Then, using these correction factors as priors, we determine the optimal correction factor for each MOSFIRE mask, making the reasonable assumption that independent measurements of the strongest emission lines in individual galaxies’ spectra should be the same; i.e., the adopted mask correction factors should minimize the scatter between the line flux measurements for the same object on different masks. For galaxies observed on a single mask, these mask corrections are used to account for slit losses, but if a galaxy has been observed on more than one mask, the line fluxes from the weighted average spectrum are corrected to match the weighted average of the corrected line fluxes for individual observations. Overall, the median correction factor is with a median uncertainty of 4%, and the uncertainties on the slit-loss corrections are later included in the final line flux uncertainties for the MOSFIRE data.
@@ 2.4 JWST/NIRSpec Spectroscopy
The JWST/NIRSpec observations including mask design and strategy are described in A. L. Strom et al. (2023). The data were taken using a single configuration of the micro-shutter assembly (MSA), at an aperture position angle of 19.3 deg. The MSA shutters have an open area of approximately and are arranged in a fixed grid, so we opted to use the “midpoint” centering criterion to jointly maximize the sample size and signal-to-noise (S/N) for individual galaxies. Given these constraints, the optimal MSA design included 34 sources. Each object’s slitlet was composed of at least 3 shutters, with additional shutters included to elongate slits in cases where there were not other open shutters in the same MSA row. Additional shutters were opened to sample the background in areas with no contamination from other sources based on deep HST images.
@@
Initial reductions of the CECILIA NIRSpec data have been described in A. L. Strom et al. (2023) and N. S. J. Rogers et al. (2024), but we take this opportunity to discuss the general procedures applied to the full sample and to highlight improvements from these initial reductions. The raw, uncalibrated data files are processed through Level 1 of calwebb v1.15.1 with crds_context 1251.pmap (H. Bushouse et al., 2024) using grizli v1.11.11 (G. Brammer, 2023a). This step performs initial data quality flagging, corrections for linearity and persistence, and cosmic ray (or ”snowball”) masking to eliminate jumps in the science groups. Once accounted for, a ramp is fit to the science groups to obtain the calibrated rate files. While recent versions of calwebb now implement a 1/f noise correction, we have elected to apply NSClean (B. J. Rauscher, 2024) directly to the rate files to remove this noise from the NIRSpec data. We first mask areas of the rate files illuminated by science or background light and pixels where residual artifacts could affect the fit to the low-frequency 1/f noise. The default NSClean parameters were found to introduce high-frequency noise to the data, so we update the critical frequency, kill width, and Gaussian smoothing standard deviation to fc1/2048, kwfc/4, and buffer_sigma3, respectively, which reduces the RMS noise in the unmasked rate files for both NRS1 and NRS2.
@@
The cleaned files are further processed through portions of the Level 2 calwebb pipeline using msaexp v0.8.5 (G. Brammer, 2022). For each group of exposures: a bias term is removed from the data, slits are identified using the slit meta data file, WCS is assigned to the rate file and individual slits, and the photometric calibration step is run. We adopt the calwebb pathloss solution for a uniformly illuminated slit. This procedure corrects for the flux loss at higher wavelengths due to the increasing full width at half maximum (FWHM) of the point spread function (PSF). The uniform solution may not be appropriate for the most compact targets or objects that are observed away from slit center. However, some of these issues are mitigated by performing a spectrophotometric correction to the 1D spectra, as described later in this section.
@@
The slit cutouts are saved in individual files to be combined and processed later. These phot files have bar shadow artifacts introduced by the bar structure between the each micro-shutter in MSA. Bar shadows can be corrected using calwebb or msaexp, although in practice we find that these solutions did not work well for the deep CECILIA data and left residual noise that complicated the measurement of a global background as well as extended source emission. Instead, we follow an iterative procedure in which we create a fiducial 1D background model similar to the approach described by A. L. Strom et al. (2023), but we also create a separate model for each detector and grating combination. We then compare the spatial variations in the background emission relative to this 1D background model among all the individual object spectra to construct a custom 2D bar shadow model. Next, we evaluate the bar shadow model on the spatial-wavelength grid for each object and divide the phot file by this model. Finally, we estimate the off-source background from each bar shadow-corrected phot file and apply an additive low-order polynomial correction to match the object-specific background to the fiducial global background solution. During this stage, we also identify and mask portions of individual exposures that are affected by open or closed shutters in the MSA. A comparison of this global background approach and the commonly-adopted local/nod background method is provided in Appendix A.
@@
The depth of the CECILIA data have revealed a number of pixels with data quality issues that are unflagged in the bad pixel masks available in crds_context 1251.pmap, and the default hot pixel algorithm tends to flag and remove intense emission lines when combining the science exposures. Instead, for each science observation we use all dither positions to create a median image, effectively removing the continuum and line emission from the target. This results in an image where bad pixels are easy to flag and remove from individual 2D spectra with a simple outlier detection algorithm. When applied to the CECILIA observations, this approach successfully removes unflagged or residual bad pixels without removing object emission. With the background, bar shadows, and bad pixels removed from the phot files, we combine the science observations and extract 1D spectra using the slit_combine package in msaexp. This program determines the object trace and profile, performs a rectification of the 2D data, and extracts the 1D science spectrum in a manner similar to K. Horne (1986). The width of the modeled extraction profile increases with wavelength to account for the increasing size of the PSF.
@@
Finally, given the potential deviations from a uniform pathloss solution and general uncertainties with the flux calibration of the NIRSpec MSA, we apply a flux correction to the 1D spectrum of each target. To do so, we take the ratio of observed continuum in NIRSpec and the SED continuum predicted from the BAGPIPES fitting for each galaxy, using a rolling median that excludes emission lines and residual data artifacts. This ratio is fit with a quadratic polynomial as a function of wavelength, then multiplied to the 1D NIRSpec data to correct the observed continuum to the model continuum. Similar approaches have been taken for other NIRSpec MSA programs, such as CEERS (P. Arrabal Haro et al., 2023), Blue Jay (S. Belli et al., 2024), and EXCELS (T. M. Stanton et al., 2025), and we apply this technique to all sources with significant continuum detections (for an analysis of the faint CECILIA galaxies, see M. Raptis et al. (2025).
@@ 2.5 Line Measurements
Emission lines in the NIRSpec data are modeled using Gaussian profiles and the SED continua discussed in §2.2. We fit the emission lines in spectral windows (up to 13 in G235M and 6 in G395M) within which we assume all emission lines have the same Gaussian FWHM and redshift. While the observed continuum generally matches the SED continuum due to the application of the spectrophotometric correction, there are instances where faint emission lines are over- or under-fit if the SED continuum is used instead of a local continuum. To capture the shape of the local continuum, we also fit a residual linear function to the difference between the observed and SED continuum and subtract this from the spectrum. The total model parameters within each window, therefore, are the local continuum slope and intercept, Gaussian FWHM, redshift, and the amplitude of each Gaussian. We subtract the SED continuum from the spectrum and use the astropy (Astropy Collaboration et al., 2013, 2018, 2022) modeling package to perform a least squares fit and determine the model parameters.
This general approach is further simplified using the expected ratios of certain metal CELs. For example, the [N II]6550,85 emission lines originate from the 1D2 level, such that the [N II]6585/[N II]6550 is fixed at 2.94 from the atomic data of C. Froese Fischer & G. Tachiev (2004) and S. S. Tayal (2011). We fix the [O III] and [N II] line flux ratios to the theoretical ratios, decreasing the number of free parameters. This matches the approach required for the MOSFIRE observations of the CECILIA galaxies, as the strong line detections in the ground-based spectra are generally lower S/N and require additional model constraints to provide robust fits to the rest-optical CELs. The [S II]6718,33 and [O II]3727,29 CELs are density sensitive, but the ratio of these emission lines vary within an expected range based on the atomic data. For NIRSpec and MOSFIRE we further constrain the model parameters by restricting the range of [S II]6718/[S II]6733 and [O II]3729/[O II]3727, respectively, to the ratios predicted for an electron gas at 1ne105 cm-3.
@@
The MOSFIRE data are fit using the same astropy method with a few notable exceptions. The MOSFIRE observations show an increased level of noise, especially at wavelengths that coincide with atmospheric OH emission lines. Therefore, we do not fit local deviations from the stellar continuum and take a different approach for the error estimation. Using the reliable error spectrum measured from the MOSFIRE data, we bootstrap resample the MOSFIRE spectra 500 times and refit the emission lines using the same astropy least squares approach. The standard deviation of the resulting flux distribution is taken as the uncertainty on the MOSFIRE line fluxes. We compare these line fluxes and uncertainties from the astropy modeling to the same quantities measured from direct integration of the isolated emission lines, and we find good agreement between the two methods. We also adopt a more constrained fit to the [O II]3727,29 lines in band, as this portion of the NIR spectrum exhibits the most significant sky emission and the [O II] lines can be unresolved depending on the spatial extent of the galaxy. To reliably fit the [O II] doublet, we leave the redshift of the line doublet as a free parameter and use the FWHM measured from the [O III] lines in band to infer the FWHM of the [O II] lines in band, accounting for the change in instrumental resolution between the two bands. We perform a visual inspection of all [O II]3727,29 doublets, and in instances where both lines are resolved and uncontaminated by sky emission (fBM40, MD43, MD41, and C31) we leave the FWHM of the [O II] lines as a free parameter.

### v2

@@ 2.2 SED Fitting
We determine the spectral energy distribution (SED) for each CECILIA target using the archival photometry described above. Note that the Palomar/WIRC photometry in is generally of worse quality than other bands, but in the case of non-detections in the IRAC channels, it is required to constrain the SED redward of the Balmer break. As a result, we only include the photometric point in the fit if the galaxy is undetected in IRAC Ch 1, 2, and 3.
@@
We use BAGPIPES (carnall2018) for SED fitting, using similar methods to those described by korhonen-cuestas2025. We employ the Bayesian nested sampler Nautilus (lange2023) and the BPASS v.2.2.1 (stanway2018) stellar grids assuming a kroupa2001-like IMF with upper mass slope of 2.35 and a maximum stellar mass of 100 . Cloudy (ferland2017) grids are used to model the nebular continuum. We do not include nebular emission lines in the SED model, as the photometry has been corrected for the contributions of emission lines using the measured JWST/NIRSpec equivalent widths as well as previously-measured emission line fluxes from MOSFIRE where available. We adopt uniform priors on the dust extinction, , logarithm of the ionization parameter, stellar metallicity, and logarithm of the stellar mass, allowing each parameter to vary between , , , and respectively. Gas phase metallicity is fixed to the stellar metallicity.111Note that the SED fits are insensitive to gas-phase metallicity because the emission lines are subtracted from the photometry and are not fit with the BAGPIPES modeling. Dust extinction is modeled assuming the gordon2003 SMC dust reddening curve. The adopted priors are based on the stellar population synthesis results reported in steidel2016 for a stacked UV spectrum of KBSS galaxies; however, as discussed in §4.1, some of the CECILIA galaxies have gas-phase O/H greater than the upper limit permitted by the metallicity prior. When repeating the SED fitting with [0.05,1.0], we find that stellar masses and SFRs are unchanged. While the median stellar metallicity increases by 0.15 dex, the typical uncertainty on is 0.37 dex. Since we do not analyze the stellar metallicities here, we adopt the narrower prior on stellar metallicity.
@@
The SFH is fit using a non-parametric, continuity SFH, as described in leja2019. We define time bins of 0, 10, 100, 250, 500, 1000, 1500, 2000, 2500, 3000, and 3500 Myr before the time of observation and fit for the change in SFR between each bin. In practice, the last time bin is sometimes excluded by the redshift of the galaxy when older than the age of the Universe at that epoch. The prior for changes in SFR between adjacent bins is set to the BAGPIPES default for a continuity SFH; i.e., a Student’s t distribution centered at 0 with a scale and 2 degrees of freedom. Such a prior is weighted against dramatic changes in SFR between time bins, generally producing smoother SFHs. Notably, the nominal SFR reported by BAGPIPES is measured on 100 Myr timescales by default, which produces an apparent maximum specific star formation rate (sSFR) of and artificial linear features to the distribution of objects in the SFR-M∗ plane. However, the underlying SFH used for the SED fitting and total mass estimation includes the full set of time bins described above. The resulting distribution of SED-fit SFRs and stellar masses for the CECILIA sample and their parent population are discussed further in Sec. 2.8.
@@ 2.3 Keck/MOSFIRE Spectroscopy
The Keck/MOSFIRE observations of KBSS galaxies have been described extensively in past work (steidel2014; strom2017). In brief, KBSS used MOSFIRE’s configurable slit unit (CSU) to observe galaxies at a time in a single atmospheric window; individual galaxies have observations spanning the , , and bands, which collectively cover Å for our highest-priority targets. Individual mask configurations were designed with -wide slits, resulting in a spectral resolution of . Observations were acquired using a two-position dither sequence along the slit, using the standard dither spacing of . Individual integration times were 180 s in band and 120 s in the and bands; total integration times range from hr/band, with a typical total integration time of hr/band. In general, once a spectroscopic redshift was determined, galaxies continued to be observed until the strongest emission lines (including [O II], H, [O III], H, and [N II]) were detected at , so some of the fainter galaxies in the sample have substantially longer exposure times.
@@
1D spectra are extracted using MOSPEC,333https://github.com/allisonstrom/mospec an interactive IDL-based analysis tool developed for faint emission line spectroscopy with MOSFIRE. We adopt a Gaussian weight profile and optimally extract the spectra using the approach from horne1986. Examples of 1D MOSFIRE data of KBSS galaxies can be found in steidel2014. To account for slit losses, which can differ from band to band depending on observing conditions, we use the Markov Chain Monte Carlo (MCMC) method from strom2017. First, we compare observations of stars that are observed on each MOSFIRE mask with their NIR magnitudes. Then, using these correction factors as priors, we determine the optimal correction factor for each MOSFIRE mask, making the reasonable assumption that independent measurements of the strongest emission lines in individual galaxies’ spectra should be the same; i.e., the adopted mask correction factors should minimize the scatter between the line flux measurements for the same object on different masks. For galaxies observed on a single mask, these mask corrections are used to account for slit losses, but if a galaxy has been observed on more than one mask, the line fluxes from the weighted average spectrum are corrected to match the weighted average of the corrected line fluxes for individual observations. Overall, the median correction factor is with a median uncertainty of 4%, and the uncertainties on the slit-loss corrections are later included in the final line flux uncertainties for the MOSFIRE data.
@@ 2.4 JWST/NIRSpec Spectroscopy
The JWST/NIRSpec observations including mask design and strategy are described in strom2023. The data were taken using a single configuration of the micro-shutter assembly (MSA), at an aperture position angle of 19.3 deg. The MSA shutters have an open area of approximately and are arranged in a fixed grid, so we opted to use the “midpoint” centering criterion to jointly maximize the sample size and signal-to-noise (S/N) for individual galaxies. Given these constraints, the optimal MSA design included 34 sources. Each object’s slitlet was composed of at least 3 shutters, with additional shutters included to elongate slits in cases where there were not other open shutters in the same MSA row. Additional shutters were opened to sample the background in areas with no contamination from other sources based on deep HST images.
@@
Initial reductions of the CECILIA NIRSpec data have been described in strom2023 and rogers2024, but we take this opportunity to discuss the general procedures applied to the full sample and to highlight improvements from these initial reductions. The raw, uncalibrated data files are processed through Level 1 of calwebb v1.15.1 with crds_context 1251.pmap (calwebb_v1.15.1) using grizli v1.11.11 (grizli_v1.9.11). This step performs initial data quality flagging, corrections for linearity and persistence, and cosmic ray (or ”snowball”) masking to eliminate jumps in the science groups. Once accounted for, a ramp is fit to the science groups to obtain the calibrated rate files. While recent versions of calwebb now implement a 1/f noise correction, we have elected to apply NSClean (rauscher2024) directly to the rate files to remove this noise from the NIRSpec data. We first mask areas of the rate files illuminated by science or background light and pixels where residual artifacts could affect the fit to the low-frequency 1/f noise. The default NSClean parameters were found to introduce high-frequency noise to the data, so we update the critical frequency, kill width, and Gaussian smoothing standard deviation to fc1/2048, kwfc/4, and buffer_sigma3, respectively, which reduces the RMS noise in the unmasked rate files for both NRS1 and NRS2.
@@
The cleaned files are further processed through portions of the Level 2 calwebb pipeline using msaexp v0.8.5 (msaexp). For each group of exposures: a bias term is removed from the data, slits are identified using the slit meta data file, WCS is assigned to the rate file and individual slits, and the photometric calibration step is run. We adopt the calwebb pathloss solution for a uniformly illuminated slit. This procedure corrects for the flux loss at higher wavelengths due to the increasing full width at half maximum (FWHM) of the point spread function (PSF). The uniform solution may not be appropriate for the most compact targets or objects that are observed away from slit center. However, some of these issues are mitigated by performing a spectrophotometric correction to the 1D spectra, as described later in this section.
@@
The slit cutouts are saved in individual files to be combined and processed later. These phot files have bar shadow artifacts introduced by the bar structure between the each micro-shutter in MSA. Bar shadows can be corrected using calwebb or msaexp, although in practice we find that these solutions did not work well for the deep CECILIA data and left residual noise that complicated the measurement of a global background as well as extended source emission. Instead, we follow an iterative procedure in which we create a fiducial 1D background model similar to the approach described by strom2023, but we also create a separate model for each detector and grating combination. We then compare the spatial variations in the background emission relative to this 1D background model among all the individual object spectra to construct a custom 2D bar shadow model. Next, we evaluate the bar shadow model on the spatial-wavelength grid for each object and divide the phot file by this model. Finally, we estimate the off-source background from each bar shadow-corrected phot file and apply an additive low-order polynomial correction to match the object-specific background to the fiducial global background solution. During this stage, we also identify and mask portions of individual exposures that are affected by open or closed shutters in the MSA. A comparison of this global background approach and the commonly-adopted local/nod background method is provided in Appendix A.
@@
The depth of the CECILIA data have revealed a number of pixels with data quality issues that are unflagged in the bad pixel masks available in crds_context 1251.pmap, and the default hot pixel algorithm tends to flag and remove intense emission lines when combining the science exposures. Instead, for each science observation we use all dither positions to create a median image, effectively removing the continuum and line emission from the target. This results in an image where bad pixels are easy to flag and remove from individual 2D spectra with a simple outlier detection algorithm. When applied to the CECILIA observations, this approach successfully removes unflagged or residual bad pixels without removing object emission. With the background, bar shadows, and bad pixels removed from the phot files, we combine the science observations and extract 1D spectra using the slit_combine package in msaexp. This program determines the object trace and profile, performs a rectification of the 2D data, and extracts the 1D science spectrum in a manner similar to horne1986. The width of the modeled extraction profile increases with wavelength to account for the increasing size of the PSF.
@@
Finally, given the potential deviations from a uniform pathloss solution and general uncertainties with the flux calibration of the NIRSpec MSA, we apply a flux correction to the 1D spectrum of each target. To do so, we take the ratio of observed continuum in NIRSpec and the SED continuum predicted from the BAGPIPES fitting for each galaxy, using a rolling median that excludes emission lines and residual data artifacts. This ratio is fit with a quadratic polynomial as a function of wavelength, then multiplied to the 1D NIRSpec data to correct the observed continuum to the model continuum. Similar approaches have been taken for other NIRSpec MSA programs, such as CEERS (arrabal2023), Blue Jay (belli2024), and EXCELS (stanton2025), and we apply this technique to all sources with significant continuum detections (for an analysis of the continuum-faint CECILIA galaxies, see raptis2025.
@@ 2.5 Line Measurements
Emission lines in the NIRSpec data are modeled using Gaussian profiles and the SED continua discussed in §2.2. We fit the emission lines in spectral windows (up to 13 in G235M and 6 in G395M) within which we assume all emission lines have the same Gaussian FWHM and redshift. While the observed continuum generally matches the SED continuum due to the application of the spectrophotometric correction, there are instances where faint emission lines are over- or under-fit if the SED continuum is used instead of a local continuum. To capture the shape of the local continuum, we also fit a residual linear function to the difference between the observed and SED continuum and subtract this from the spectrum. The total model parameters within each window, therefore, are the local continuum slope and intercept, Gaussian FWHM, redshift, and the amplitude of each Gaussian. We subtract the SED continuum from the spectrum and use the astropy (astropy2013; astropy2018; astropy2022) modeling package to perform a least squares fit and determine the model parameters.
This general approach is further simplified using the expected ratios of certain metal CELs. For example, the [N II]6550,85 emission lines originate from the 1D2 level, such that the [N II]6585/[N II]6550 is fixed at 2.94 from the atomic data of froesefischer2004 and tayal2011. We fix the [O III] and [N II] line flux ratios to the theoretical ratios, decreasing the number of free parameters. This matches the approach required for the MOSFIRE observations of the CECILIA galaxies, as the strong line detections in the ground-based spectra are generally lower S/N and require additional model constraints to provide robust fits to the rest-optical CELs. The [S II]6718,33 and [O II]3727,29 CELs are density sensitive, but the ratio of these emission lines vary within an expected range based on the atomic data. For NIRSpec and MOSFIRE we further constrain the model parameters by restricting the range of [S II]6718/[S II]6733 and [O II]3729/[O II]3727, respectively, to the ratios predicted for an electron gas at 1ne105 cm-3.
@@
The MOSFIRE data are fit using the same astropy method with a few notable exceptions. The MOSFIRE observations show an increased level of noise, especially at wavelengths that coincide with atmospheric OH emission lines. Therefore, we do not fit local deviations from the stellar continuum and take a different approach for the error estimation. Using the reliable error spectrum measured from the MOSFIRE data, we bootstrap resample the MOSFIRE spectra 500 times and refit the emission lines using the same astropy least squares approach. The standard deviation of the resulting flux distribution is taken as the uncertainty on the MOSFIRE line fluxes. We compare these line fluxes and uncertainties from the astropy modeling to the same quantities measured from direct integration of the isolated emission lines, and we find that the percent difference between the resulting line fluxes and uncertainties from the two methods are 1% and 5%, respectively. We also adopt a more constrained fit to the [O II]3727,29 lines in band, as this portion of the NIR spectrum exhibits the most significant sky emission and the [O II] lines can be unresolved depending on the spatial extent of the galaxy. To reliably fit the [O II] doublet, we leave the redshift of the line doublet as a free parameter and use the FWHM measured from the [O III] lines in band to infer the FWHM of the [O II] lines in band, accounting for the change in instrumental resolution between the two bands. We perform a visual inspection of all [O II]3727,29 doublets, and in instances where both lines are resolved and uncontaminated by sky emission (fBM40, MD43, MD41, and C31) we leave the FWHM of the [O II] lines as a free parameter.